2 3 The ammeter reads l=348 A. Find ii. (V=8.3 V, Ri=214 Ω, R2=4.61 Ω, R3-2.49 Ω.) .09 A Submit Answer Incorrect. Tries 3/7 Previous Tries Find 12 Submit Answer Tries 0/7 Find E Submit Answer Tries 0/7
Answer #1

a ) Writing the loop equation for left loop we have

8.3 - I1*2.14 - 3.48*2.49 = 0

I1 = -0.17 A

b) I1 + I2 = I

-0.17 + I2 = 3.48

I2 = 3.65 A

c) Writing loop equation for right loop we have

3.48*2.49 + 3.65*4.61 - E = 0

E = 25.5 V